### v1

# DB-Explore: Automated Database Exploration and Instruction Synthesis for Text-to-SQL

Abstract
Recent text-to-SQL systems powered by large language models (LLMs) have demonstrated remarkable performance in translating natural language queries into SQL. However, these systems often struggle with complex database structures and domain-specific queries, as they primarily focus on enhancing logical reasoning and SQL syntax while overlooking the critical need for comprehensive database understanding. To address this limitation, we propose DB-Explore, a novel framework that systematically aligns LLMs with database knowledge through automated exploration and instruction synthesis. DB-Explore constructs database graphs to capture complex relational schemas, leverages GPT-4 to systematically mine structural patterns and semantic knowledge, and synthesizes instructions to distill this knowledge for efficient fine-tuning of LLMs. Our framework enables comprehensive database understanding through diverse sampling strategies and automated instruction generation, bridging the gap between database structures and language models. Experiments conducted on the SPIDER and BIRD benchmarks validate the effectiveness of DB-Explore, achieving an execution accuracy of 52.1% on BIRD and 84.0% on SPIDER. Notably, our open-source implementation, based on the Qwen2.5-coder-7B model, outperforms multiple GPT-4-driven text-to-SQL systems in comparative evaluations, and achieves near state-of-the-art performance with minimal computational cost.
DB-Explore: Automated Database Exploration and Instruction Synthesis for Text-to-SQL
Haoyuan Ma1, Yongliang Shen1, Hengwei Liu1, Wenqi Zhang1, Haolei Xu1, Qiuying Peng2 Jun Wang2, Weiming Lu1 1Zhejiang University, 2 OPPO Research Institute {mahaoyuan, syl, luwm}@zju.edu.cn {pengqiuying, wangjun7}@oppo.com
footnotetext: Corresponding author.
1 Introduction
Text-to-SQL systems have emerged as a crucial bridge between humans and structured data, enabling natural language interaction with databases without requiring SQL expertise. While recent advances in large language models (LLMs), including ChatGPT (Liu et al., 2023), GPT-4 (Achiam et al., 2023), and Qwen2 (Hui et al., 2024), have significantly improved these systems’ capabilities, demonstrating enhanced comprehension of user queries and the ability to generate complex SQL queries, they still face fundamental challenges in understanding complex database structures and generating accurate SQL queries, particularly for domain-specific applications (Zhang et al., 2024; Liu et al., 2023).
A key limitation of current approaches is their reliance on direct query translation without deep understanding of the underlying database structure and semantics. Existing methods primarily focus improving prompting strategies (Dong et al., 2023; Rajkumar et al., 2022a; Liu et al., 2023; Sun et al., 2023a) or enhancing the reasoning ability of LLMs through supervised fine-tuning (Li et al., 2023a; Scholak et al., 2021; Wang et al., 2020), but often overlook the critical role of database comprehension in query generation. While techniques like schema linking (Wang et al., 2020; Li et al., 2023a), question decomposition (Pourreza and Rafiei, 2023) and self-correction (Wang et al., 2023) have attempted to address this gap, they typically treat the database as a static structure rather than a rich source of semantic knowledge that can inform the query generation process.
We argue that effective Text-to-SQL systems must first develop a comprehensive understanding of the database through systematic exploration before attempting query generation. This exploration should capture not only the structural relationships between tables but also the semantic patterns and domain-specific knowledge embedded within the schema. However, automated database exploration presents several challenges: (1) Complex structural relationships: Real-world databases often contain intricate networks of table relationships that are difficult to capture and represent effectively. (2) Hidden semantic patterns: Domain-specific knowledge embedded in schema elements (table names, column names, etc.) is often implicit and requires careful analysis to uncover. (3) Query complexity progression: Understanding how simple queries can be systematically extended to handle more complex scenarios requires a structured approach to knowledge acquisition.
To address these challenges, we propose DB-Explore, a novel framework that transforms Text-to-SQL training through systematic database exploration and instruction synthesis. At the core of our approach is a graph-based database representation technique called DB Graph, which captures both structural relationships and semantic patterns within the database. This representation serves as a foundation for three key components: (1) Semantic Knowledge Extraction: We develop techniques to uncover domain-specific knowledge embedded in schema elements, enabling more contextually appropriate query generation. (2) Structural Pattern Mining: By analyzing the DB Graph, our system identifies common join patterns, table relationships, and query templates that reflect the database’s architectural design. (3) Progressive Instruction Synthesis: Using the extracted knowledge, we automatically generate training examples that progress from simple to complex queries, allowing models to gradually master increasingly sophisticated query patterns.
Our framework leverages these components to create comprehensive training datasets that capture both the structural and semantic complexity of the target database. This approach fundamentally differs from existing methods by treating database exploration as a crucial prerequisite for effective query generation, rather than relying solely on input-output pairs or schema information. The main contributions of this work are:
  1. 1.
    We introduce DB-Explore, a framework that revolutionizes Text-to-SQL training by incorporating systematic database exploration and knowledge extraction.
  2. 2.
    We develop DB Graph, a novel representation that captures both structural relationships and semantic patterns within databases, enabling more effective exploration and learning.
  3. 3.
    We present techniques for automated instruction synthesis that leverage extracted database knowledge to create progressively complex training examples.
  4. 4.
    We demonstrate significant improvements in query generation accuracy and complexity handling across multiple benchmark datasets with minimal computational cost.
2 Related Work
Figure 1: Overall framework of DB-Explore. Our framework operates through three principal phases: (1) Database Graph Construction, (2) Database Multi-Knowledge Exploration: semantic knowledge extraction, structural pattern mining, progressive instruction synthesis, and filtering mechanisms, and (3) Supervised Fine-Tuning executing database-adaptive model training. The synthesis phase systematically produces database-consistent training corpora through multi-stratum knowledge alignment.
Early work in the field of Text-to-SQL generation primarily relied on rule-based methods to parse natural language queries and map them to valid SQL statements (Li and Jagadish, 2014; Popescu et al., 2003, 2004). More recently, LLMs such as GPT-4 (Achiam et al., 2023) have facilitated this task by leveraging their powerful text priors to generate SQL (Li et al., 2023b). Building upon this foundation, diverse technical approaches have emerged, such as RESD-SQL, DTS-SQL and RSL-SQL (Li et al., 2023a; Pourreza and Rafiei, 2024; Cao et al., 2024). These methods focus on enhancing schema selection and employ schema linking techniques to improve the identification of database schemas. Another line of research addresses complex query decomposition—DIN-SQL, PTD-SQL and QPL (Pourreza and Rafiei, 2023; Luo et al., 2024; Eyal et al., 2023) adopt stepwise problem decomposition to break down complex SQL generation into manageable subtasks. Candidate generation strategies emerge in approaches like MSC-SQL, SQL-PaLM, Distillery and XiYan (Gorti et al., 2024; Sun et al., 2023a; Maamari et al., 2024; Gao et al., 2024b), which boost accuracy through multi-candidate SQL generation and selection. DAIL-SQL (Gao et al., 2024a) introduces a skeleton-aware few-shot demonstration selection to enhance relevance of examples. MAC (Wang et al., 2023) and EPI-SQL (Liu and Tan, 2024) incorporate self-correction frameworks for SQL refinement.
Notably, recent open-source 7B/13B models (e.g., CodeLlama (Roziere et al., 2023), Qwen2 (Hui et al., 2024)) demonstrate competitive Text-to-SQL performance. These technical advances have catalyzed the emergence of multiple pre-training/fine-tuning frameworks leveraging large-scale synthetic data. CODES (Li et al., 2024a)) incrementally pre-trains the StarCoder (Li et al., 2023c) model using a corpus related to SQL and incorporates various external plugins to enhance Text-to-SQL performance. DTS-SQL (Pourreza and Rafiei, 2024) employs SFT on a dedicated schema linking model to reduce the input complexity of the model and improve performance. SENSE (Yang et al., 2024b) enhances the model’s preference on correct SQL by synthesizing strong data and weak data and direct preference optimization (DPO) on LLMs (Rafailov et al., 2024). ROUTE (Qin et al., 2024) unifies multi-task fine-tuning through task-specific data synthesis. However, these methodologies predominantly focus on enhancing LLMs’ direct SQL generation capabilities while overlooking comprehensive database understanding.
Our work aims to enhance the Text-to-SQL capabilities of LLMs by dynamic mining the knowledge underlying database structure and semantics. Through database exploration and data synthesis, we extract database knowledge and inject it into base model, reducing potential risks caused by insufficient knowledge of databases.
3 Methodology
In this section, we present a detailed explanation of DB-Explore, a fully automated framework that leverages LLMs to explore databases, generate data, and perform supervised fine-tuning. As illustrated in Figure 1, the framework comprises three core components: 1) Database Graph Construction, 2) Database Multi-Knowledge Exploration, and 3) Supervised Fine-Tuning.
DB-Explore initiates with the construction of DB Graph, which captures the intricate topological relationships within the database. By applying various sampling strategies, we extract subgraphs containing diverse information. DB-Explore then converts this database knowledge into a instruction dataset through semantic knowledge extraction, structural pattern mining, and progressive instruction synthesis. To ensure the quality of the generated instructions, an instruction filtering mechanism is employed. Finally, the extracted knowledge is injected into the base model via supervised fine-tuning. In the following subsections, we present a detailed description of each component of DB-Explore.
3.1 Database Graph Construction and Sampling
To effectively model both structural and semantic relationships within databases, we formalize database schemas as DB graph, as illustrated in Step 1 of Figure 1. The graph comprises nodes representing database columns and edges encoding relational constraints. Two edge types are defined: Intra-table edges connect columns within the same table, enabling direct access without JOIN operations; Inter-table edges represent foreign key relationships across tables, bridging subgraphs into a unified structure. Edge weights are assigned based on co-occurrence frequency of connected columns in seed queries, prioritizing frequently accessed schema patterns that reflect real-world query behaviors.
To enable more effective exploration and learning, we propose two complementary sampling strategies:
  1. 1.
    Random Walk Sampling: Starting from selected nodes, the walker iteratively traverses adjacent nodes via randomly chosen edges until reaching a predefined sampling size. This method ensures comprehensive schema coverage but may overlook high-priority relationships.
  2. 2.
    Empirical Weighted Sampling: Normalized edge weights guide traversal probabilities, biasing selection toward high-weight edges. This strategy emphasizes schema subgraphs aligned with user query patterns, enhancing instruction relevance.
Both strategies generate subgraphs as input for subsequent database multi-knowledge exploration, balancing semantic diversity via random walks and user intent alignment via weighted sampling.
3.2 Database Multi-Knowledge Exploration
We propose a three-pronged exploration and data synthesis framework to explore database by instruction synthesis while minimizing computational overhead: 1) Semantic Knowledge Extraction, 2) Structural Pattern Mining, and 3) Progressive Instruction Synthesis. The prompt templates shown in Figure 1 can be found in Appendix A.
3.2.1 Semantic Knowledge Extraction
Semantic knowledge extraction mechanism leverages LLMs’ internal knowledge to discover latent schema relationships through a self-guided paradigm. The process initiates by sampling schema subgraphs $S$ from the DB Graph via random walk sampling, containing table/column metadata and structural relationships.
Self-instruction has been shown to be an effective way to synthesize data (Wang et al., 2022). The instruction synthesis pipeline operates through a self-instruct framework initialized using manually-crafted template. We prompt the LLM with $S$ and randomly sampled seed instructions as demonstrations to synthesize novel queries $Q$ through schema-constrained synthesis. For instance, given columns "order_amount" and "customer_country", the model generates: "List customers from Canada with total orders exceeding 500" The generated ⟨$Q$, $S$⟩ pairs are subsequently feedback to the LLM to produce executable SQL responses $R$.
To achieve optimal semantic diversity while preserving logical validity, we implement temperature scaling ($\tau$=0.7) during generation. This technique yields paraphrased variants such as "Find Canadian clients spending over 500" while maintaining schema constraints. The iterative process progressively enriches the instruction pool through random sampling, thereby expanding coverage of potential user intents.
3.2.2 Structural Pattern Mining
Unlike semantic knowledge extraction, structural pattern mining is designed to generate schema-grounded SQL queries through explicit modeling of relational topologies derived from the DB Graph. It prioritizes foreign-key dependencies and latent connectivity patterns observed in seed instructions, while strategically balancing single-table operations with multi-tables queries through weighted edge sampling. We propose a SQL-first generation paradigm to ensure structural accuracy by initially producing executable SQL statements that rigorously adhere to schema constraints. These syntactically valid queries are subsequently translated into natural language instructions via inverse parsing, effectively reducing hallucination risks compared to open-ended synthesis.
As Figure 2 showed we first employ empirical weighted sampling to select schema subgraphs $S$, prioritizing high-frequency foreign-key edges "price.ID $=$ production.id". Then for each $S$, SQL statements $R$ are automatically generated by traversing all nodes using:
  • JOIN: Link tables via foreign keys
  • Filters: Derived from column data types
  • Aggregations: Applied to numerical columns
A representative SQL output for subgraph $S$:
"SELECT price.id FROM price INNER JOIN production ON price.id = production.id WHERE production.country = ’USA’ AND price.price < 15000"
Surviving SQL is then translated into natural language queries via back instruct (Shen et al., 2023), where the LLM rephrases R into natural language while preserving structural accuracy converting the above SQL to:
"List IDs of North American-manufactured vehicles under \$15,000"
Figure 2: Structural Pattern Mining Illustration.
3.2.3 Progressive Instruction Synthesis
In this section, we introduce a progressive instruction synthesis mechanism that progressively enriches query complexity to address the challenge of generating complex queries from simple user inputs. As illustrated in Figure 3, SQL queries are formalized as syntax trees where the root represents the core SELECT statement and leaf nodes correspond to schema-specific constraints. Each leaf node maps to a modular condition that can be independently translated into natural language, establishing a bidirectional bridge between SQL semantics and user intent.
Figure 3: Progressive Instruction Synthesis Illustration.
The evolution process initiates with a base query sampled from the DB Graph. Through iterative refinement: 1. Condition Augmentation: Schema-derived nodes (foreign keys, constraints) are recursively appended to the syntax tree via DB Graph-guided sampling. 2. Complexity Scaling: The LLM orchestrates tree expansion by injecting compound conditions such as nested WHERE clauses and multi-table JOIN while preserving syntactic validity. 3. Paraphrase Generation: Each evolved SQL tree is verbalized into diverse natural language expressions through template-guided LLM rewriting.
This progressive paradigm generates instruction pairs spanning atomic queries to multi-hop reasoning tasks, effectively simulating real-world query evolutionary pattern.
3.3 Instruction Filtering
Progressive instruction synthesis often generates a significant amount of noisy data due to hallucinations of LLMs. To mitigate noise data, we implement a dual-stage validation pipeline:
  1. 1.
    Execution Verification:
    All generated SQL candidates R are executed against the target database. Queries are discarded based on three criteria:
    • Syntax errors
    • Invalid column/table references
    • Execution time exceeding 25 seconds
  2. 2.
    Consistency Verification:
[…]
    • Faithful translation between natural language intent and SQL logic
    • Adherence to schema constraints in S
[…]
3.4 Supervised Fine-Tuning and Inference
To inject the database knowledge into base model, we implement a two-stage supervised fine-tuning framework using the synthesized instruction dataset $\textlangle Q,D,R,S\textrangle$. Addressing the challenge of schema complexity versus limited context windows, user queries undergo schema linking to extract critical schema elements $S$ before SQL generation. This cascaded approach reduces irrelevant schema noise and minimizes token consumption.
To enhance the model’s comprehension of database structures during inference, we integrates a Value Retrieval Module which implements a coarse-to-fine matching mechanism via BM25 indexing and Longest Common Subsequence (LCS) algorithms (Li et al., 2024a), resolving token mismatches. Additionally, metadata augmentation is utilized, embedding database descriptions, column types, foreign keys, and domain knowledge into prompts.
Schema Linking Model
Schema linking model trained to identify schema $S_{i}$ relevant to user queries $Q_{i}$. Given above $T={\{(Q_{i},D_{i})\}}$ , wherein the input prompt and target response generated is relevant schema $S_{i}$. The primary objective during supervised fine-tuning for schema linking model is to minimize the following loss function:
$Loss=-\frac{1}{|T|}\sum_{i=1}^{|T|}p(S_{i}|Q_{i},D_{i},M_{i})$ (1)
where $p$ represents the probability assigned to $S_{i}$ by the schema linking model. Integrating, $M_{i}$ represents the metadata of $D_{i}$ and $Q_{i}$.
SQL Generation Model
Conditioned on the schema subset $S_{i}$ from the linking model, this component maps $\textlangle Q_{i},S_{i}\textrangle$to executable SQL $R_{i}$. We then fine-tune the LLM while minimizing the following loss function:
$Loss=-\frac{1}{|T|}\sum_{i=1}^{|T|}p(R_{i}|Q_{i},S_{i},M_{i})$ (2)
Here $M_{i}$ represents the metadata of $S_{i}$ and $Q_{i}$.
4 Experiments
Methods Spider Bird
Dev-EX Dev-TS Dev-EX Dev-VES
Prompting with Closed-Source LLMs
MCS-SQL + GPT-4 (Lee et al., 2025) 89.5 - 63.4 64.8
XIYAN (Gao et al., 2024b) 89.7 - 73.3 -
CHASE-SQL + Gemini-1.5 (Pourreza et al., 2024) 87.6 - 73.1 -
GPT-4 (Achiam et al., 2023) 72.9 64.9 46.4 49.8
DIN-SQL + GPT-4 (Pourreza and Rafiei, 2023) 82.8 74.2 50.7 58.8
DAIL-SQL + GPT-4 (Gao et al., 2024a) 83.5 76.2 54.8 56.1
MAC-SQL + GPT-4 (Wang et al., 2023) 86.8 82.8 59.4 66.2
TA-SQL + GPT-4 (Qu et al., 2024) 85.0 - 56.2 -
MAG-SQL + GPT-4 (Xie et al., 2024) 85.3 - 61.1 -
Fine-Tuning with Open-Source LLMs 7B
DTS-SQL + DeepSeek-7B (Pourreza and Rafiei, 2024) 82.7 78.4 55.8 60.3
CODES-7B (Li et al., 2024a) 85.4 80.3 57.2 58.8
SENSE-7B (Yang et al., 2024b) 83.2 81.7 51.8 -
ROUTE + Qwen2.5-7B (Qin et al., 2024) 83.6 77.5 55.9 57.4
Ours: DB-Explore + Qwen2.5-Coder-7B 84.0 79.3 52.1 55.8
Table 1: Experimental results on SPIDER and BIRD benchmarks. Asterisks ’’ denote performance re-evaluated through the official open-source implementation by ROUTE. Within the Fine-Tuning with Open-Source LLMs 7B group, top-performing entries are bold with runner-up scores underlined.
Benchmarks
(1) SPIDER The Spider dataset (Yu et al., 2018) is a large-scale, cross-domain benchmark for Text-to-SQL tasks, designed to evaluate the generalization ability of models across diverse database structures. It consists of 8,659 training samples, and 1,034 development samples, spanning 200 databases and 138 domains. Spider provides a robust challenge for models, emphasizing adaptability to unfamiliar database schemas. (2) BIRD The BIRD dataset (Li et al., 2023b) is a more challenging Text-to-SQL benchmark compared to SPIDER, featuring 95 real-world databases across 37 professional domains, with a total storage size of 33.4GB. BIRD emphasizes the challenges of handling large-scale, noisy, and incomplete database values, alongside the need for efficient SQL query generation
Evaluation Metrics
In our experiments, we use three key evaluation metrics to assess the performance of Text-to-SQL models. Execution Accuracy (EX) (Yu et al., 2018) measures the proportion of queries where the predicted SQL query produces the exact output as the ground truth, serving as a direct indicator of correctness. Test-Suite Accuracy (TS) (Zhong et al., 2020) stands out as a more trustworthy metric than EX. For BIRD, we report EX along with the Valid Efficiency Score (VES) (Li et al., 2023b), which is used to evaluate the execution efficiency of accurately generated SQL queries. Unlike EX, in VES the score for an accurately generated SQL query is no longer fixed at 1. Instead, it is determined by the ratio of the execution time of the ground truth SQL query to that of the predicted SQL query.
Implementation Details
[…]
Baselines
In our experimental evaluation, we conduct comprehensive comparisons with state-of-the-art Text-to-SQL approaches across multiple technical paradigms. Our baselines can be categorized into two groups: 1) The prompting methods with closed-source LLMs MCS-SQL (Lee et al., 2025) and Chase-SQL (Guo et al., 2021), which enhance SQL generation accuracy through multi-prompt engineering and systematic comparison of candidate results. MAC-SQL (Wang et al., 2023) and MAG-SQL (Xie et al., 2024) employ multi-agent architectures for error correction and consensus building. TA-SQL (Qu et al., 2024) implement task alignment mechanisms to address LLMs’ hallucination. DIN-SQL (Pourreza and Rafiei, 2023) adopts task decomposition to handle complex queries through step-wise problem simplification, while DAIL-SQL (Gao et al., 2024a) selects high quality demonstration through skeleton-aware few-shot demonstration selection. 2) Fine-tuning with open-source LLMs 7B DTS-SQL (Pourreza and Rafiei, 2024) utilizes a multistage strategies based methods with two-stage fine-tuning and generation pipeline. SENSE (Yang et al., 2024b) introduces preference learning between strong and weak data in the Text-to-SQL domain. Codes (Li et al., 2024a) enhances its Text-to-SQL capabilities by incorporating a large amount of pretraining data and external plugins. On the other hand, ROUTE (Qin et al., 2024) introduces multi-task learning into the Text-to-SQL domain.
5 Results and Analysis
5.1 Main Results
Methods Spider
Dev-EX Dev-TS
Ours: DB-Explore + Qwen2.5-Coder-7B 84.0 79.3
- w/o Semantic Knowledge Extraction 82.9 78.4
- w/o Structural Pattern Mining 82.1 77.9
- w/o Progressive Instruction Synthesis 83.1 78.7
Table 2: Ablation study on DB-Explore, complete method of DB-Explore is bold
Results on SPIDER
Table 1 presents the performance comparison between our method and baseline approaches on the SPIDER development set. The results demonstrate that prompt-based baselines achieve promising performance when augmented with GPT-4, while methods such as SENSE, CODES, and ROUTE bridge the performance gap between open-source models and GPT-4 through extensive data synthesis. In contrast, our database exploration mechanism effectively enhances open-source models performance with minimal computational cost. On the SPIDER benchmark, our approach achieves 84.0% EX, surpassing all 7B-parameter model-based methods except CODES, and outperforming certain GPT-4 based approaches such as DAIL-SQL and DIN-SQL.
Figure 4: Data volume analysis of different data synthesis methods on SPIDER, CODES only reports the storage size for pre-training data.
Results on BIRD
Our method achieves 52.1% EX with Qwen2.5-Coder-7B on BIRD development set, as shown in Table 1. However, in practice, our approach is not fully applicable to the BIRD dataset, as the discrepancy between training and evaluation database. Nevertheless, the data synthesis paradigm demonstrates foundational Text-to-SQL enhancement capabilities, achieving this performance with only 40MB augmented training samples. We obtained a result of 55.8% on the VES metric of the BIRD dataset. The VES metric effectively reflects the execution efficiency of SQL, highlighting the high efficiency of our data synthesis method. As our approach focuses on improving LLMs’ understanding of databases rather than handling complex logic, the synthesized data maintains a high degree of consistency with the database structure. This reduces redundant generation and unnecessary database operations, leading to significant progress on the VES metric.
Data Analysis
Unlike existing approaches that focus on enhancing LLMs’ generalization capabilities for SQL generation, DB-Explore adopts a database-centric data synthesis paradigm that prioritizes mining database-specific knowledge. This fundamental distinction enables our method to achieve competitive performance with significantly reduced data requirements. As illustrated in Figure 4, our approach achieves the lowest computational cost among all data synthesis-based Text-to-SQL methods. Specifically, DB-Explore demonstrates superior data efficiency by attaining higher EX scores on the SPIDER benchmark than SENSE and ROUTE while using only 79% and 21% of their respective data volumes. Although our performance on SPIDER slightly trails CODES (84.0% vs. 85.4%), this comparison merits particular attention given that our supervised fine-tuning data constitutes merely 0.15% of the pre-training data required by CODES.
5.2 Ablation Study
Study on Semantic Knowledge Extraction
We conduct ablation experiments on the semantic knowledge extraction component of DB-Explore, which aligns LLMs with database domains through schema-aware semantic extraction for precise SQL generation and diversified instruction synthesis. As shown in Table 2, removing this component causes a 1.1% performance drop on the SPIDER benchmark, confirming its necessity for maintaining model efficacy.
Study on Structural Pattern Mining
As evidenced in Table 2, the structural pattern mining component is crucial to DB-Explore, as it explicitly models relational database schemas. This module enhances the LLMs’ ability to understand database topologies and structural dependencies, distinguishing it from baseline methods focused solely on direct SQL generation. Removing structural pattern mining results in a 1.9% performance drop on the SPIDER benchmark, primarily due to foreign-key errors and column selection mismatches. Notably, structural pattern mining is closely linked to empirical weighted sampling who ensures that the synthesized data better aligns with the distribution of the seed data.
Study on Progressive Instruction Synthesis
Progressive instruction synthesis enhances the complexity of instruction data by iteratively adding conditions to existing instructions. This framework effectively addresses the complexity limitations of LLM-based data synthesis. In our experiments, initial data synthesis performed well for 2-table and 3-table instructions, but for instructions involving more than three tables, LLMs exhibited a high error rate. The resulting limitations in instruction complexity significantly restrict the SQL generation capability of LLMs. As shown in Table 2, progressive instruction synthesis yielded a 0.9% performance improvement for DB-Explore. However, during the experiment, progressive instruction synthesis led to a shift in the synthesized data distribution, which resulted in suboptimal performance on the SPIDER dataset. Nonetheless, we believe that high-quality complex data can further elevate LLMs’ Text-to-SQL capabilities in real-world applications.
6 Conclusion
In this paper, we propose DB-Explore, a framework for database exploration, instruction synthesis, and fine-tuning that enhances the LLMs’ understanding of databases by extracting dynamic database knowledge, thereby unlocking the potential of LLMs in Text-to-SQL tasks. Our approach, through database multi-knowledge exploration, effectively minimizes the risks associated with insufficient database knowledge in SQL generation. We demonstrate the efficacy and superiority of our method on recent LLMs across several benchmark tests. The results show that our approach achieves satisfactory performance in Text-to-SQL with minimal computational cost. In the future, we aim to explore more database knowledge, larger LLMs, and more efficient fine-tuning frameworks to achieve robust and efficient Text-to-SQL solutions.
Limitations
While our exploration demonstrates promising results, several important limitations warrant acknowledgment. First, the framework introduces modest computational cost and GPT-4o API consumption. Although these costs remain substantially lower than comparable methods, we encourage future work to explore more efficient data generation and fine-tuning paradigms. Second, our upper-bound analysis reveals a persistent performance gap between our approach and inference frameworks directly leveraging closed-source models. Finally, due to computational constraints, the framework’s effectiveness on larger language models (e.g., Qwen2-72B (Hui et al., 2024)) remains unverified, leaving its scalability to state-of-the-art architectures as open research questions.
Appendix A Appendix
Our DB-Explore framework is designed with various prompt templates aimed at enhancing GPT-4o’s performance in tasks such as semantic knowledge extraction, structural pattern mining, and progressive instruction synthesis. We generated a total of 6,497 question-SQL pairs on the SPIDER dataset and 6,687 question-SQL pairs on the BIRD dataset. In this appendix, we provide the prompts of Database Multi-Knowledge Exploration used in our pipeline as shown in Figure 5 to 7.
Figure 5: The prompt of Semantic Knowledge Extraction.
Figure 6: The prompt of Structural Pattern Mining.
Figure 7: The prompt of Progressive Instruction Synthesis.